APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Sunday June 18, 2006 (MJD 53905) ---=== OBSERVING TEAM ===--- Mike Harvanek Olena Malanushenko Shannon Watters ---=== OBSERVING PLAN ===--- Blind you with Science! ---=== OBSERVING SUMMARY ===--- The sky was marginally photometric at the beginning of the night but there were no imaging strips available so we started with spectroscopy. We finished plate 1723 and then switched to imaging. We tried to image strip 290 S but had problems with the pointing/skippy. By the time we straightened those problems out the seeing was bad (>2.5) so we switched back to spectroscopy. We had to reboot the crate for ad4. While on plate 2539 the horrible seeing (>3") caused problems acquiring all the guide stars so we didn't get an exposure finished until the plate was expiring. There wasn't enough signal so we had to declare it bad. We were able to finish 1 more plate, 2182, before we closed at 08:56Z due to a combination of humidity (>50%) and integrated dust-hours (>20,000). After closing Russet checked the outside dust count with a portable meter and found the level was substantionally lower than the tower count, so we opened. We should have checked this on our own before closing--observer error. We opened up and started on plate 2537 but SOS gave us errors for both exposures (frame 39846 & 39847) for the r1 camera: "ABORT: Reject science: 443 saturated rows" "ABORT: Unable to reduce science exposure" We declared the exposures bad. Could Brian Yanny please look at the data and let us know if a change in observing strategy is needed. While closing we started a manual fill of the spectro dewar when French called and told us to put it into "Auto" mode. It was photometric all night except from ~5:00Z to ~7:00Z. We don't have an integrated dust count because of innaccurate readings from the tower meters. The mirror probably doesn't need cleaning but please check it. ---=== OBSERVING LOG ===--- Afternoon: ---------- Spectro collimation was checked but not adjusted. No problems to report. Night: ------ 02:25Z As instructed, we manually filled the spectro dewars until they were spitting LN2 out of their relief valves. This took 54 lbs. When we switched the controller back to the "auto" position to stop the fill it kept filling. We had to set it to the "off" position to stop the fill. 02:55Z Cart 1, plt 1723, tile 1229, chunk87. No GSOGTF; used FK5. Plate finished (exposures: 900, 900, 800). 03:15Z Howard suggested that we reboot the DIMM after its failure last night. It was rebooted and is working at the moment. 04:45Z Switched to imaging. 05:24Z While trying to image Strip 290 S (run 6216) we couldn't get a nu error from Skippy until frame 42. The nu error was > 18" so we made an adjustment and the nu error was fine. We billed this time lost due to pointing as "Telescope hardware problem". 05:40Z We backed up and began another drift on 290 S (run 6217) when the watcher complained that there was no response on ad4 (frame 21). The ad4 monitor wasn't scrolling so we rebooted its crate and it came back. During all of this the wind speed increased and shifted from SW to SE, the dust count climbed to ~3500, the sky became marginally photometric, and the seeing became horrible (>3"). We billed this lost time to seeing (Other weather problem). 06:00Z We switched back to spectroscopy. 06:10Z Cart 3, plt 2539, tile 8209, segue146. No GSOGTF; FK5 used. We found 3 guide stars but not the others. We played with the focus and the scale, cleared the offsets, went to an FK5 again, and finally found guide stars but the seeing was still terrible. We continued billing this lost time to seeing. After finding all the guide stars and tweaking the focus and scale the seeing was down to 2.5". 06:42Z Precals finally started. 06:49Z Began 1200 sec. exposure. This was the only exposure on this plate because it expired at 0707Z. The (S/N)^2 was less than 2 for 3 of the cameras so the exposure was declared "BAD". 07:17Z Cart 9, plt 2182, tile 8054, segue117. No GSOGTF; used offset. Plate finished (exposures: 1200, 900, 900, 900). 08:50Z The DIMM worked properly until we shut it down because of dust. 08:56Z Closed the enclosure because of a combination of humidity (>50%) and an integrated dust count of 20,000 dust-hours. 09:35Z Russet checked the outdoor dust count with a handheld dust count meter and got counts substantially lower (~1700) than the tower reading which was ~5000. There must be a problem with the tower counter so we opened and were on the sky again by 09:44Z. The time lost was billed as "High dust". 09:44Z Cart 2, plt 2537, tile 8207, segue146. No GSOGTF; used FK5. (exposures: 900, 900). After the first exposure (frame 39846) SOS gave us "ABORT: Reject Science: 443 saturated rows" and "ABORT: Unable to reduce science exposure" on camera r1. Another exposure was being taken so we let it finish and the SOS results were the same. We declared the exposures "BAD". endNight rejoiced. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 6215 00:14Z 00:34Z 100 O -20.07 -15.21 44 ignore 6216 04:53Z 05:24Z 290 S 93.55 101.53 67 ignore 6217 05:29Z 05:40Z 290 S 94.39 97.36 33 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el 6216 46 4 0 0.000 0.000 0.000 0.000 -0.00000 0 0 6216 56 4 50 0.500 0.169 0.200 0.157 0.00463 105 53 6217 14 4 99 1.400 0.187 0.200 0.188 0.00369 117 56 ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ 6216 20 1 50 71 80 1.580 0.324 1.447 0.306 -0.00487 6216 20 1 50 71 67 1.987 0.260 1.427 0.306 -0.00480 6216 20 1 50 71 74 2.427 0.224 1.540 0.424 -0.00518 6216 20 1 50 71 74 2.623 0.164 1.358 0.392 -0.00456 6216 20 1 50 71 69 2.878 0.221 1.585 0.260 -0.00533 6216 33 1 51 73 64 1.502 0.240 -0.896 0.290 0.00301 6216 33 1 51 73 60 2.021 0.266 -0.863 0.418 0.00290 6216 33 1 51 73 62 2.372 0.297 -0.619 0.379 0.00208 6216 33 1 51 73 68 2.598 0.293 -0.551 0.375 0.00185 6216 33 1 51 73 74 2.906 0.262 -0.761 0.264 0.00256 ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): YES QA Procedures Done (y/n): YES UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 39825, plate -9999 ------- 00:13 39825 0.0 bias ----- sequence 39826, plate 1723 ------- 00:36 39826 10.0 flat 00:38 39827 2.0 arc ----- sequence 39828, plate -9999 ------- 01:46 39828 0.0 bias 01:51 39829 0.0 bias ----- sequence 39830, plate 1723 ------- 15.4 15.4 19.8 21.9 DONE 03:32 39830 10.0 flat 03:34 39831 2.0 arc 03:54 39832 900.1 target 04:11 39833 900.1 target 04:28 39834 800.1 target ----- sequence 39835, plate 2539 ------- 1.9 3.2 0.5 1.3 BAD 06:44 39835 10.0 flat 06:46 39836 2.0 arc 07:12 39837 1200.1 target ----- sequence 39838, plate 2182 ------- 19.5 19.9 17.0 29.7 DONE 07:25 39838 10.0 flat 07:28 39839 2.0 arc 07:51 39840 1200.1 target 08:08 39841 900.1 target 08:37 39842 900.1 target 08:55 39843 900.1 target ----- sequence 39844, plate 2537 ------- Unable to reduce. BAD 09:57 39844 10.0 flat 09:59 39845 2.0 arc 10:22 39846 900.1 target 10:40 39847 900.1 target ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 03:17Z 1 1723 45.36 0.0001 58.06 0.0066 217.08 0.0021 1.000320 05:14Z imager 73.24 0.0000 51.90 0.0000 106.03 0.0000 1.000000 06:43Z 3 2539 215.82 -0.0009 50.88 0.0031 69.05 0.0097 1.000300 07:34Z 9 2182 -67.85 -0.0065 76.69 0.0027 118.94 0.0243 1.000260 10:07Z 2 2537 168.46 0.0033 51.57 0.0020 -28.66 0.0094 0.999990 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL Stays: JL ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 03:17Z 1 1723 1.00032 -3160 -1608 -150 45 58.1 20.3 7 194 --- 1.2 05:14Z imager 1.00000 0 -546 -457 73 51.9 19.3 11 180 06:43Z 3 2539 1.00030 -2961 -1074 200 216 50.9 15.9 24 143 --- 2.5 07:34Z 9 2182 1.00026 -2568 -1053 -50 -68 76.6 15.7 22 134 --- 1.9 10:07Z 2 2537 0.99999 99 1258 0 168 51.6 15.9 14 142 --- 2.3 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 23:03Z 74 30 14 265 (W) 1225 - clear 23:37Z 77 33 16 256 (WSW) 1336 - clear 00:10Z 76 31 14 266 (W) 1259 - clear 00:41Z 73 29 6 262 (W) 1265 - clear 01:11Z 74 28 11 265 (W) 1107 - clear 01:42Z 71 26 6 265 (W) 1119 - clear 02:06Z 69 24 3 260 (W) 1210 - clear 02:38Z 67 22 5 198 (SSW) 1131 - clear 03:10Z 68 23 5 219 (SW) 1157 - clear 03:41Z 67 23 9 208 (SSW) 1222 - clear 04:11Z 68 23 11 193 (SSW) 1405 - clear 04:41Z 67 23 10 178 (S) 1802 - clear 05:11Z 66 22 9 181 (S) 1739 - clear 05:41Z 62 33 21 136 (SE) 3093 - clear 06:12Z 61 37 16 139 (SE) 1706 - clear 06:43Z 60 40 24 142 (SE) 2849 - clear 07:14Z 60 41 21 138 (SE) 4266 - clear 07:44Z 60 41 22 135 (SE) 2931 - clear 08:15Z 60 41 19 152 (SSE) 6747 - clear 08:45Z 60 42 22 149 (SSE) 6791 - clear 09:17Z 60 41 20 150 (SSE) 5482 - clear 09:48Z 61 38 17 153 (SSE) 4935 - clear 10:20Z 61 36 12 157 (SSE) 4294 - clear 10:52Z 61 32 8 173 (S) 3850 - clear ---=== TELESCOPE STATUS ===--- 01:19Z Enclosure doors opened, fans and blowers on, louvers opened. 02:25Z Enclosure off. Manually filled spectro dewars. They took 54 lbs. 08:56Z Enclosure on, doors closed, fans and blowers off, louvers closed. 09:44Z Enclosure off, doors open. 10:50Z Enclosure on, doors closed. Telescope stowed at 30 degrees with cartridge 2 (plate 2537) mounted. Counterweights at 250. Spectro auto-fill connected. Spectro dewar 11:05Z 108.10 lb, 25.46 psi Imager dewar 11:05Z 156,60 lb, 18.53 psi No interlocks bypassed. ---=== SOFTWARE USED ===--- IOP/SOP: v4_24_0b Watcher: v2_34_0 MCP: v6_3_2 TPM: tpm_v3_3_0 AstroDa: v15_12_1 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v2_18 SoS: v5_1_4 hoggPT: v1_6_9 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.8550 -5.8440 Axial B -5.6840 -5.6770 Axial C 0.9430 0.9420 Trans D -9.9990 -9.9940 Lateral E 10.3632 10.3759 Lateral F 11.9380 11.9380 GALILS Commanded: 4700. -2450. 3000. -7800. 3300. 3000. Actual: 4700. -2450. 3000. -7800. 3300. 3000. SETMIR VALUES PriDesOrient: 0.00 -1.68 7.05 385.25 92.02 PriOrient: 0.00 -1.52 7.02 384.99 93.81 SECONDARY: ---------- Focus: 0.00 Air Temp.: 24.5 Alt.: 30.032259 MIGS TONIGHT NOMINAL Axial A 1.5370 1.5420 Axial B 1.0570 1.0470 Axial C 1.1660 1.1650 Trans D -0.1290 -0.1240 GALILS Commanded: 1593463. 1603643. 1554295. -6800. -5900. Actual: 1593450. 1603650. 1554300. -6800. -5900. SETMIR VALUES SecDesOrient: 1257.00 15.87 0.00 -50.00 136.76 SecOrient: 1257.00 15.88 -0.01 -49.56 136.73 ---=== PROBLEMS IN DETAIL ===---